# Estimation of discrete distributions in relative entropy, and the deviations of the missing mass

### v2

@@ Abstract
We study the problem of estimating a distribution over a finite alphabet from an i.i.d. sample, with accuracy measured in relative entropy (Kullback-Leibler divergence). While optimal expected risk bounds are known, high-probability guarantees remain less well-understood. First, we analyze the classical Laplace (add-one) estimator, obtaining matching upper and lower bounds on its performance and showing its optimality among confidence-independent estimators. We then characterize the minimax-optimal high-probability risk, which is attained via a simple confidence-dependent smoothing technique. Interestingly, the optimal non-asymptotic risk exhibits an additional logarithmic factor over the ideal asymptotic risk. Next, motivated by scenarios where the alphabet exceeds the sample size, we investigate methods that adapt to the sparsity of the distribution at hand. We introduce an estimator using data-dependent smoothing, for which we establish a high-probability risk bound depending on two effective sparsity parameters. As part of the analysis, we also derive a sharp high-probability upper bound on the missing mass.
@@ 1.1 Problem setting
An important feature of the Kullback-Leibler divergence, compared to other common divergences between probability distributions such as the total variation and Hellinger distances, is that it penalizes significant underestimation of true frequencies. To consider an extreme case, if the estimator $\widehat{P}_{n}$ assigns a probability $\widehat{p}_{j}=0$ to a class $j$ with probability $p_{j}\neq 0$, then the Kullback-Leibler divergence $\mathrm{KL}(P,\widehat{P}_{n})$ is infinite. This aligns with the needs of various applications: for instance, in a forecasting context where classes correspond to different outcomes, assigning a probability of 00 to outcomes that are in fact possible would constitute a severe underestimation of the underlying uncertainty. In addition, in the context of language models, one may be interested in generating new sentences that are not present in the training corpus; this requires assigning positive probabilities to sequence of words that have not been observed.
@@
In particular, if all classes $j=1,\dots,d$ have nonzero probability, then $\overline{P}_{n}$ converges to $P$ as $n\to\infty$ at a rate of $1/\sqrt{n}$ in distribution, is asymptotically normal and efficient [vdV98, §5.2–5.6]. As a result, $2n\cdot\mathrm{KL}(P,\overline{P}_{n})$ converges in distribution to a $\chi^{2}$ distribution with $d-1$ degrees of freedom. Together with a standard tail bound on $\chi^{2}$ distributions, this implies the following guarantee: for any fixed $d\geqslant 2$, $P\in\mathcal{P}_{d}$, and $\delta\in(0,1)$, one has
@@
It should be emphasized that the lack of uniformity over the distribution $P$ of the point-wise asymptotic guarantee (3) is not merely an artifact of its formulation, but instead reflects a fundamental limitation of the estimator $\overline{P}_{n}$ itself. Indeed, the empirical distribution $\overline{P}_{n}$ assigns a probability of 00 to classes that do not appear; such a configuration may occur in a finite sample, especially in the presence of rare classes. Hence, the MLE is generally inadequate for the purpose of density estimation in relative entropy, due to its propensity to underestimate uncertainty and to produce overly sharp probability estimates.
@@ 1.2 Existing guarantees
(where we used (5) and that $\sqrt{d}\log^{5/2}d\lesssim d$). In particular, this removes a $\sqrt{\log(1/\delta)}$ factor in the deviation term compared to (9), leading to optimal guarantees in the larger range $\log(1/\delta)\lesssim d^{1/5}$. Nevertheless, the bound still deteriorates in the regime $\log(1/\delta)\gg d^{1/5}$.
@@ 1.5 Notation
Throughout this work, we let $n\geqslant 1$ denote the sample size and $d\geqslant 2$ the number of classes. If $A$ is a finite set, we denote by $|A|$ its cardinality. We identify the set of probability distributions on $[d]=\{1,\dots,d\}$ with the set of probability vectors $\mathcal{P}_{d}=\{(p_{1},\dots,p_{d})\in\mathbb{R}_{+}^{d}:\sum_{j=1}^{d}p_{j}=1\}$, where for $1\leqslant j\leqslant d$ we let $p_{j}$ denote the probability of the class $j$. For $j\in\{1,\dots,d\}$, we let $\delta_{j}\in\mathcal{P}_{d}$ denote the Dirac mass at $j$, identified with the $j$-th basis vector in $\mathbb{R}^{d}$. Given two probability distributions $P=(p_{1},\dots,p_{d})\in\mathcal{P}_{d}$ and $Q=(q_{1},\dots,q_{d})\in\mathcal{P}_{d}$, we define the Kullback-Leibler divergence or relative entropy between $P$ and $Q$ by
@@
with the convention that $p\log(p/q)$ equals 00 if $p=0$, and +$\infty$+ ∞ if $p>0$ but $q=0$. For $u,v\in\mathbb{R}^{+}$, we let $D(u,v)=u\log(\frac{u}{v})-u+v$ with similar conventions. Since $\sum_{j=1}^{d}p_{j}=\sum_{j=1}^{d}q_{j}=1$, we have
@@ 2.2 Lower bound for confidence-independent estimators
A few comments may help clarify the meaning of Theorem 2. First, one should think of the parameter $\kappa\geqslant 1$ as an absolute constant. The condition (17) states that the estimator $\widehat{P}_{n}$ achieves a bound of order $d/n$ with positive probability for any distribution $P$. This holds in particular when the estimator $\widehat{P}_{n}$ achieves an optimal in-expectation bound $\mathbb{E}_{P}[\mathrm{KL}(P,\widehat{P}_{n})]\leqslant\kappa d/n$, and more generally when the estimator achieves an optimal bound of $\kappa d/n$ in the regime of constant (bounded away from 00 and $1$) confidence level. The first condition applies with $\kappa=1$ to the Laplace estimator, in light of its in-expectation bound (5). We refer to such estimators, including those that are optimal in expectation, as “confidence-independent”.
@@ 3.2 Lower bound for confidence-dependent estimators
The idea behind Lemma 1 is quite simple, and can be summarized as follows: let $P$ be either a Dirac mass at $1$, namely $P=\delta_{1}=(1,0,\dots,0)$, or a mixture of the form $(1-\frac{\log(1/\delta)}{n})\delta_{1}+\frac{\log(1/\delta)}{n}\delta_{j}$ for some $j=2,\dots,d$. Then, regardless of which of these distributions $P$ is, with probability at least of order $\delta$, only the first class is observed ($X_{1}=\dots=X_{n}=1$). In this case, it is impossible to tell which distribution $P$ is. In order to avoid incurring a large error in the first case where $P=\delta_{1}$, one must assign a large probability to the first class, and thus a low probability to all remaining classes. However, this entails a large error in the second case. Specifically, the extra $\log(d)$ factor comes from the fact that in the second case where $P$ puts some mass on a class $j\in\{2,\dots,d\}$, no information on $j$ is available if only the first class is observed. Hence, all the remaining mass must be shared among the $d-1$ remaining classes $j=2,\dots,d$, effectively dividing by $d-1$ the per-class probability and thus inflating the relative entropy.
@@ 4.1 Minimax lower bounds for sparse distributions
However, despite the similarity in rates, there are qualitative differences between the Gaussian setting and the multinomial setting that we consider. Indeed, it follows from the results of Agrawal [Agr22] (see Lemma 4 below) that when $P\in\mathcal{P}_{s,d}$, the empirical distribution achieves an upper bound in squared Hellinger distance of order $s/n$ with probability at least $1-e^{-s}$. This rate no longer features the extra $\log(ed/s)$ factor; by contrast, the extra $\log(ed/s)$ factor does appear in the Gaussian model, even when the error is measured in squared Hellinger distance. Roughly speaking, this difference stems from the fact that in the multinomial model, unlike in the Gaussian model, the amount of “noise” in a coordinate $j=1,\dots,d$ decays when the magnitude of the corresponding coefficient parameter $p_{j}$ goes to 00.
@@
A second important feature of the lower bound of Proposition 1 (and of its consequence (25) in the non-sparse case) is that it holds with high probability. Such a statement is stronger than minimax lower bounds that are usually found in the literature, which hold either in expectation or with constant probability. For estimation of non-sparse discrete distributions, lower bounds in expectation are proven in [HJW15, KOPS15], while [CKT24] obtains a lower bound with constant probability. To appreciate the difference, consider the $d/n$ lower bound in the non-sparse case. Then, a $d/n$ lower bound in expectation is in principle compatible with the following behavior: a risk of $1$ with probability $d/n$, and a risk of $0\ll d/n$ with high probability $1-d/n$. In contrast, a stronger lower bound with constant probability rules out such a behavior: it asserts that a lower bound of order $d/n$ must hold with probability bounded away from 00, say $10\%$ or $80\%$. However, even this lower bound does not rule out the possibility that the estimator achieves an error significantly smaller than $d/n$ (say, of 00) with nontrivial probability—respectively, with probability $90\%$ or $20\%$. A high-probability lower bound such as (25) excludes such a behavior: it asserts that a lower bound of order $d/n$ holds with overwhelming probability, effectively ruling out the possibility that the estimator “gets lucky”. In addition, the probability with which the lower bound (25) holds, which behaves as $1-e^{-d/c}$, is best possible, as shown by the convergence of the properly scaled risk of the MLE to a $\chi_{d-1}^{2}$ distribution as $n\to\infty$.
Proposition 1 is, to our knowledge, the first minimax lower bound that holds with probability exponentially close to $1$. We believe that such lower bounds could be of broader interest in statistical estimation, beyond the present context of estimation of discrete distributions.
We note that the classical method for proving lower bounds in statistical estimation (put forward by Ibragimov and Has’minskii [IH81], see [Tsy09, Chapter 2] and [Wai19, Chapter 15]), which consists in a reduction from estimation to testing, followed by a lower bound for testing though Fano’s inequality, actually provides a high-probability lower bound. However, the probability estimate that the commonly used version of this inequality (see e.g. [Wai19, Proposition 5.12 p. 502]) leads to is weaker than (25). Specifically, it leads to a lower bound that holds with probability $1-c/d$ for some constant $c$, in contrast with the lower bound (25) with probability $1-e^{-d/c}$. As it turns out, in the non-sparse case, one could also obtain a probability of $1-e^{-d/c}$ by instead using the sharp version of Fano’s inequality (e.g., [Tsy09, Lemma 2.10]) and further simplifying the resulting bound. However, the reduction from estimation to testing is not straightforward in the sparse case, essentially because the Kullback-Leibler divergence does not behave like a distance over sparse distributions with distinct supports.
@@ 4.2.1 Effective sparsity parameters
Perhaps the most natural notion of “effective support size of $P$ under sample size $n$” is the typical number of distinct classes that would appear on an i.i.d. sample of size $n$ from $P$. As we will explain briefly, this leads to the following definition:
@@ 4.2.3 High-probability upper bounds for sparse distributions
For instance, the results of Section 3 (Theorems 3 and 4) imply that for $d\geqslant 5000$, $0<\delta<e^{-2}$ and $10<\varepsilon<$ 1, a sample size of
@@ 5 High-probability bound on the missing mass
We first show that $\sum_{j^{*}<j\leqslant d}p_{j}<\varepsilon$. If $j^{*}=d$, the sum equals 00 and the property holds; we thus assume that $j^{*}\leqslant d-1$. In this case, let $j^{\prime}$ denote the largest integer in $\{j^{*}+1,\dots,d\}$ such that $\sum_{j^{*}<j\leqslant j^{\prime}}p_{j}<\varepsilon$ (by definition of $j^{*}$, this inequality holds for $j^{\prime}=j^{*}+1$, hence $j^{\prime}\geqslant j^{*}+1$). We need to show that $j^{\prime}=d$. We proceed by contradiction and assume that $j^{\prime}\leqslant d-1$. In this case, one has $\sum_{j^{*}<j\leqslant j^{\prime}}p_{j}<\varepsilon$ while $\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}\geqslant\varepsilon$. But since $\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}\in\mathcal{S}(P)$, this also implies that $\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}\geqslant\overline{\varepsilon}\geqslant 2\varepsilon$, hence $p_{j^{\prime}+1}=\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}-\sum_{j^{*}<j\leqslant j^{\prime}}p_{j}>\varepsilon$. But this contradicts the fact that $p_{j^{\prime}+1}\leqslant p_{j^{*}+1}<\varepsilon$, proving the claim of this paragraph.
@@
However, there are situations in which the second term in (55) dominates and leads to a suboptimal bound. For instance, consider the case where $p_{j}\asymp 1/d$ for $21\leqslant j\leqslant d/$ 2, while $p_{j}\asymp 2^{-(j-d/2)}/d$ for $d/2<j\leqslant d$, and the regime of constant probability $\delta\in(0,1)$. For $d\log d\ll n\ll 2^{d}$, the typical missing mass is at most of order $\mathbb{E}_{P}[M_{n}]\asymp 1/n$. However, in this regime one has $d_{n}^{+}\asymp d$, hence the bound (55) is of order $\sqrt{d}/n$, which is suboptimal by a $\sqrt{d}$ factor. This $\sqrt{d}$ factor can be removed by resorting to Theorem 6.

### v3

@@ Abstract
We study the problem of estimating a distribution over a finite alphabet from an i.i.d. sample, with accuracy measured in relative entropy (Kullback-Leibler divergence). While optimal bounds on the expected risk are known, high-probability guarantees remain less well-understood. First, we analyze the classical Laplace (add-one) estimator, obtaining matching upper and lower bounds on its performance and establishing its optimality among confidence-independent estimators. We then characterize the minimax-optimal high-probability risk and show that it is achieved by a simple confidence-dependent smoothing technique. Notably, the optimal non-asymptotic risk incurs an additional logarithmic factor compared to the ideal asymptotic rate. Next, motivated by regimes in which the alphabet size exceeds the sample size, we investigate methods that adapt to the sparsity of the underlying distribution. We introduce an estimator using data-dependent smoothing, for which we establish a high-probability risk bound depending on two effective sparsity parameters. As part of our analysis, we also derive a sharp high-probability upper bound on the missing mass.
@@ 1.1 Problem setting
An important feature of the Kullback-Leibler divergence, compared to other common divergences between probability distributions such as the total variation and Hellinger distances, is that it penalizes significant underestimation of true frequencies. To consider an extreme case, if the estimator $\widehat{P}_{n}$ assigns a probability $\widehat{p}_{j}=0$ to a class $j$ with probability $p_{j}\neq 0$, then the Kullback-Leibler divergence $\mathrm{KL}(P,\widehat{P}_{n})$ is infinite. This aligns with the needs of various applications: for instance, in a forecasting context where classes correspond to different outcomes, assigning a probability of 0 to outcomes that are in fact possible would constitute a severe underestimation of the underlying uncertainty. In addition, in the context of language models, one may be interested in generating new sentences that are not present in the training corpus; this requires assigning positive probabilities to sequence of words that have not been observed.
@@
In particular, if all classes $j=1,\dots,d$ have nonzero probability, then $\overline{P}_{n}$ converges to $P$ as $n\to\infty$ at a rate of $1/\sqrt{n}$ in distribution, is asymptotically normal and efficient [vdV98, §5.2–5.6]. As a result, $2n\cdot\mathrm{KL}(P,\overline{P}_{n})$ converges in distribution to a $\chi^{2}$ distribution with d1d-1 degrees of freedom. Together with a standard tail bound on $\chi^{2}$ distributions, this implies the following guarantee: for any fixed $d\geqslant 2$, $P\in\mathcal{P}_{d}$, and $\delta\in(0,1)$, one has
@@
It should be emphasized that the lack of uniformity over the distribution $P$ of the point-wise asymptotic guarantee (3) is not merely an artifact of its formulation, but instead reflects a fundamental limitation of the estimator $\overline{P}_{n}$ itself. Indeed, the empirical distribution $\overline{P}_{n}$ assigns a probability of 0 to classes that do not appear; such a configuration may occur in a finite sample, especially in the presence of rare classes. Hence, the MLE is generally inadequate for the purpose of density estimation in relative entropy, due to its propensity to underestimate uncertainty and to produce overly sharp probability estimates.
@@ 1.2 Existing guarantees
(where we used (5) and that $\sqrt{d}\log^{5/2}d\lesssim d$). In particular, this removes a $\sqrt{\log(1/\delta)}$ factor in the deviation term compared to (9), leading to optimal guarantees in the larger range $\log(1/\delta)\lesssim d^{1/5}$. Nevertheless, the bound still deteriorates in the regime $\log(1/\delta)\gg d^{1/5}$. This being said, unlike other results discussed above, the guarantee from [CSS23] establishes concentration of the error $\mathrm{KL}(P,\widehat{P}_{n})$ around its expectation, rather than merely a deviation bound.
@@
Finally, after a first version of the present paper was disseminated, a concurrent work of van der Hoeven, Olkhovskaia and van Erven [vdHOvE25] simplified the estimator from [vdHZCB23] and refined its analysis, obtaining a guarantee of order $\{d\log\log d+\log(d)\log(1/\delta)\}/n$ which is near-optimal up to the $\log\log d$ factor.
@@ 1.5 Notation
Throughout this work, we let $n\geqslant 1$ denote the sample size and $d\geqslant 2$ the number of classes. If $A$ is a finite set, we denote by $|A|$ its cardinality. We identify the set of probability distributions on $d]=\{1,\dots,d\}$ with the set of probability vectors $\mathcal{P}_{d}=\{(p_{1},\dots,p_{d})\in\mathbb{R}_{+}^{d}:\sum_{j=1}^{d}p_{j}=1\}$, where for $1\leqslant j\leqslant d$ we let $p_{j}$ denote the probability of the class $j$. For $j\in\{1,\dots,d\}$, we let $\delta_{j}\in\mathcal{P}_{d}$ denote the Dirac mass at $j$, identified with the $j$-th basis vector in $\mathbb{R}^{d}$. Given two probability distributions $P=(p_{1},\dots,p_{d})\in\mathcal{P}_{d}$ and $Q=(q_{1},\dots,q_{d})\in\mathcal{P}_{d}$, we define the Kullback-Leibler divergence or relative entropy between $P$ and $Q$ by
@@
with the convention that $p\log(p/q)$ equals 0 if $p=0$, and $\infty$ if $p>0$ but $q=0$. For $u,v\in\mathbb{R}^{+}$, we let $D(u,v)=u\log(\frac{u}{v})-u+v$ with similar conventions. Since $\sum_{j=1}^{d}p_{j}=\sum_{j=1}^{d}q_{j}=1$, we have
@@ 2.2 Lower bound for confidence-independent estimators
A few comments may help clarify the meaning of Theorem 2. First, one should think of the parameter $\kappa\geqslant 1$ as an absolute constant. The condition (17) states that the estimator $\widehat{P}_{n}$ achieves a bound of order $d/n$ with positive probability for any distribution $P$. This holds in particular when the estimator $\widehat{P}_{n}$ achieves an optimal in-expectation bound $\mathbb{E}_{P}[\mathrm{KL}(P,\widehat{P}_{n})]\leqslant\kappa d/n$, and more generally when the estimator achieves an optimal bound of $\kappa d/n$ in the regime of constant (bounded away from 0 and $1$) confidence level. The first condition applies with $\kappa=1$ to the Laplace estimator, in light of its in-expectation bound (5). We refer to such estimators, including those that are optimal in expectation, as “confidence-independent”.
@@ 3.2 Lower bound for confidence-dependent estimators
The idea behind Lemma 1 is quite simple, and can be summarized as follows: let $P$ be either a Dirac mass at $1$, namely $P=\delta_{1}=(1,0,\dots,0)$, or a mixture of the form $(1-\frac{\log(1/\delta)}{n})\delta_{1}+\frac{\log(1/\delta)}{n}\delta_{j}$ for some $j=2,\dots,d$. Then, regardless of which of these distributions $P$ is, with probability at least of order $\delta$, only the first class is observed ($X_{1}=\dots=X_{n}=1$). In this case, it is impossible to tell which distribution $P$ is. In order to avoid incurring a large error in the first case where $P=\delta_{1}$, one must assign a large probability to the first class, and thus a low probability to all remaining classes. However, this entails a large error in the second case. Specifically, the extra $\log(d)$ factor comes from the fact that in the second case where $P$ puts some mass on a class $j\in\{2,\dots,d\}$, no information on $j$ is available if only the first class is observed. Hence, all the remaining mass must be shared among the d1d-1 remaining classes $j=2,\dots,d$, effectively dividing by d1d-1 the per-class probability and thus inflating the relative entropy.
@@ 4.1 Minimax lower bounds for sparse distributions
However, despite the similarity in rates, there are qualitative differences between the Gaussian setting and the multinomial setting that we consider. Indeed, it follows from the results of Agrawal [Agr22] (see Lemma 4 below) that when $P\in\mathcal{P}_{s,d}$, the empirical distribution achieves an upper bound in squared Hellinger distance of order $s/n$ with probability at least $1-e^{-s}$. This rate no longer features the extra $\log(ed/s)$ factor; by contrast, the extra $\log(ed/s)$ factor does appear in the Gaussian model, even when the error is measured in squared Hellinger distance. Roughly speaking, this difference stems from the fact that in the multinomial model, unlike in the Gaussian model, the amount of “noise” in a coordinate $j=1,\dots,d$ decays when the magnitude of the corresponding coefficient parameter $p_{j}$ goes to 0.
@@
A second important feature of the lower bound of Proposition 1 (and of its consequence (25) in the non-sparse case) is that it holds with high probability. Such a statement is stronger than minimax lower bounds that are usually found in the literature, which hold either in expectation or with constant probability. For estimation of non-sparse discrete distributions, lower bounds in expectation are proven in [HJW15, KOPS15], while [CKT24] obtains a lower bound with constant probability. To appreciate the difference, consider the $d/n$ lower bound in the non-sparse case. Then, a $d/n$ lower bound in expectation is in principle compatible with the following behavior: a risk of $1$ with probability $d/n$, and a risk of $0\ll d/n$ with high probability 1d/n1-d/n. In contrast, a stronger lower bound with constant probability rules out such a behavior: it asserts that a lower bound of order $d/n$ must hold with probability bounded away from 0, say $10\%$ or $80\%$. However, even this lower bound does not rule out the possibility that the estimator achieves an error significantly smaller than $d/n$ (say, of 0) with nontrivial probability—respectively, with probability $90\%$ or $20\%$. A high-probability lower bound such as (25) excludes such a behavior: it asserts that a lower bound of order $d/n$ holds with overwhelming probability, effectively ruling out the possibility that the estimator “gets lucky”. In addition, the probability with which the lower bound (25) holds, which behaves as $1-e^{-d/c}$, is best possible, as shown by the convergence of the properly scaled risk of the MLE to a $\chi_{d-1}^{2}$ distribution as $n\to\infty$.
We note that the classical method for proving lower bounds in statistical estimation (put forward by Ibragimov and Has’minskii [IH81], see [Tsy09, Chapter 2] and [Wai19, Chapter 15]), which consists in a reduction from estimation to testing, followed by a lower bound for testing though Fano’s inequality, actually provides a high-probability lower bound. However, the probability estimate that the commonly used version of this inequality (see e.g. [Wai19, Proposition 5.12 p. 502]) leads to is weaker than (25). Specifically, it leads to a lower bound that holds with probability 1c/d1-c/d for some constant $c$, in contrast with the lower bound (25) with probability $1-e^{-d/c}$. As it turns out, in the non-sparse case, one could also obtain a probability of $1-e^{-d/c}$ by instead using the sharp version of Fano’s inequality (e.g., [Tsy09, Lemma 2.10]) and further simplifying the resulting bound. However, the reduction from estimation to testing is not straightforward in the sparse case, essentially because the Kullback-Leibler divergence does not behave like a distance over sparse distributions with distinct supports.
@@
We believe that minimax lower bounds that hold with probability exponentially close to $1$ (such as Proposition 1) have broad relevance in statistical estimation beyond the present setting of discrete distributions, as they provide very precise information. Although such lower bounds appear to be scarce in the literature, a recent example is [RHJR24, Theorem 2.1] which is established via the $\chi^{2}$ mutual information using tools from [CGZ16].
@@ 4.2.1 Effective sparsity parameters
Perhaps the most natural notion of “effective support size of $P$ under sample size nn” is the typical number of distinct classes that would appear on an i.i.d. sample of size $n$ from $P$. As we will explain briefly, this leads to the following definition:
@@ 4.2.3 High-probability upper bounds for sparse distributions
For instance, the results of Section 3 (Theorems 3 and 4) imply that for $d\geqslant 5000$, $0<\delta<e^{-2}$ and $0<\varepsilon<1$, a sample size of
@@ 5 High-probability bound on the missing mass
We first show that $\sum_{j^{*}<j\leqslant d}p_{j}<\varepsilon$. If $j^{*}=d$, the sum equals 0 and the property holds; we thus assume that $j^{*}\leqslant d-1$. In this case, let $j^{\prime}$ denote the largest integer in $\{j^{*}+1,\dots,d\}$ such that $\sum_{j^{*}<j\leqslant j^{\prime}}p_{j}<\varepsilon$ (by definition of $j^{*}$, this inequality holds for $j^{\prime}=j^{*}+1$, hence $j^{\prime}\geqslant j^{*}+1$). We need to show that $j^{\prime}=d$. We proceed by contradiction and assume that $j^{\prime}\leqslant d-1$. In this case, one has $\sum_{j^{*}<j\leqslant j^{\prime}}p_{j}<\varepsilon$ while $\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}\geqslant\varepsilon$. But since $\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}\in\mathcal{S}(P)$, this also implies that $\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}\geqslant\overline{\varepsilon}\geqslant 2\varepsilon$, hence $p_{j^{\prime}+1}=\sum_{j^{*}<j\leqslant j^{\prime}+1}p_{j}-\sum_{j^{*}<j\leqslant j^{\prime}}p_{j}>\varepsilon$. But this contradicts the fact that $p_{j^{\prime}+1}\leqslant p_{j^{*}+1}<\varepsilon$, proving the claim of this paragraph.
@@
However, there are situations in which the second term in (55) dominates and leads to a suboptimal bound. For instance, consider the case where $p_{j}\asymp 1/d$ for $1\leqslant j\leqslant d/2$, while $p_{j}\asymp 2^{-(j-d/2)}/d$ for $d/2<j\leqslant d$, and the regime of constant probability $\delta\in(0,1)$. For $d\log d\ll n\ll 2^{d}$, the typical missing mass is at most of order $\mathbb{E}_{P}[M_{n}]\asymp 1/n$. However, in this regime one has $d_{n}^{+}\asymp d$, hence the bound (55) is of order $\sqrt{d}/n$, which is suboptimal by a $\sqrt{d}$ factor. This $\sqrt{d}$ factor can be removed by resorting to Theorem 6.